OpenStudy (anonymous):

find the equation of the line containing the given pair of points. (-3,0),(0,4) y= ?

OpenStudy (anonymous):

Use the point to find the slope then use the point slope equation I described in your last question to get the equation

OpenStudy (anonymous):

(y2 - y1) / (x2 - x1) = m

OpenStudy (abmon98):

m=0+3/4-0=3/4 y-y1=m(x-x1) y-0=3/4(x+3) 4y=3x+12 y=3/4+3

OpenStudy (abmon98):

sorry y=3/4x+3